OpenStudy (anonymous):

Solve the equation of log (2x + 4) = log (3x -12).

OpenStudy (anonymous):

x=16

OpenStudy (anonymous):

thanks :) how did you solve that?

OpenStudy (anonymous):

I use an app. It's called Mathway helps out a lot

OpenStudy (anonymous):

oh :D

OpenStudy (triciaal):

log (2x + 4) = log (3x -12) the value of a number does not change when expressed differently example 5*2 = 10 and 6 + 4 = 10 means 5*2 = 6 + 4 10 is still 10 so for this problem .2x + 4 = 3x - 12 solve for x variable to one side divide by coefficient of variable keep the balance of the equation any questions?

OpenStudy (anonymous):

thanks for showing me how to do it.

OpenStudy (triciaal):

you are welcome!